OpenStudy (anonymous):

What is the difference between 2/3 and the repeating decimal 0.393939... ?

myininaya (myininaya):

if \[x=.\overline{39}\] then \[100x=39.\overline{39}\] so \[100x-x=39\] =>\[99x=39\] =>\[x=\frac{39}{99}=\frac{13}{33}\]

myininaya (myininaya):

the difference is 2/3-13/33=22/33-13/33=9/33=3/11

OpenStudy (anonymous):

kk thanx

myininaya (myininaya):

np

OpenStudy (curry):

zarkon take a look at my question please its so hard